OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a15 = -53, a16 = -5

OpenStudy (anonymous):

what is the difference d between two terms in the sequence ?

OpenStudy (anonymous):

48?

OpenStudy (anonymous):

correct now the formula for the nth term in arithmetic sequence is: an = a1 + (n-1)d can you find a1 ?

OpenStudy (anonymous):

I know its -725 i think but is that it?

OpenStudy (anonymous):

so lets check, lets plug a16 (or a15) into the formula: a16 = a1 + (16-1)*48 a1 = a16 - 15 * 48 = -725 so it is correct

OpenStudy (anonymous):

so now plug a1 and d into the formula for an and done :)

OpenStudy (anonymous):

so would the equation be an=-725 +48(n-1)?

OpenStudy (anonymous):

yes ! you can always verify it by plugging n=15 or 16 and check if you get the given values a15 = -725 + 48 * 14 = -53 so it is correct indeed

OpenStudy (anonymous):

(this is a good check to do in a test for example.. you can see if your answer is good or not)

OpenStudy (anonymous):

Thank you so so much! You have NO idea how much you have helped me

OpenStudy (anonymous):

yw :)